OpenStudy (anonymous):

At a high school, the probability that a student takes a science class and a history class is 0.48. The probability that a student takes a science class is 0.82, and the probability that the student takes a history class is 0.64. What is the probability (rounded to the nearest hundredth) that a student takes a history class given that the student is taking a science class? 0.39 0.75 0.30 0.59

OpenStudy (b87lar):

start with the formula that a prob of something (say, A) given something else (say, B) is \[P(A|B)=\frac{P(A,B)}{P(B)}\]

OpenStudy (b87lar):

can you identify what is A and B in the wording?

OpenStudy (anonymous):

Not really. which is which?

OpenStudy (b87lar):

P(A|B) --> P(takes history | taking science) , so A corresponds to taking history, B corresponds to taking science

OpenStudy (b87lar):

P(A,B) is the probability that a student is taking *both*

OpenStudy (b87lar):

P(B) is the prob that a student is taking science (irregardless of whether she takes history or not)

OpenStudy (b87lar):

now you can plug the numbers into the formula

OpenStudy (b87lar):

I am getting 0.58537 which, rounded up, looks like (d)

OpenStudy (anonymous):

How did you get there? If you don't mind me asking

OpenStudy (b87lar):

0.48/0.82

OpenStudy (b87lar):

0.48 is the P(A,B) and 0.82 is P(B)

OpenStudy (b87lar):

does it make sense?

OpenStudy (anonymous):

Yes that makes sense

OpenStudy (anonymous):

OK thank you! I have one more. Would you mind helping me with it?

OpenStudy (b87lar):

yes

OpenStudy (b87lar):

i mean go ahead

OpenStudy (anonymous):

The probability that you roll a 3 on a six sided die is 1/6. The probability that you flip a coin that lands on tails is 1/2. The probability that you roll a 3 on a six-sided die and you flip a coin that lands on tails is 1/12. What is the probability of flipping a coin and it landing on tails, given that you rolled a 3 on a six-sided die? Are these two events independent?

OpenStudy (anonymous):

The answer choices: P(T, 3) = 1/12; therefore, events are independent because P(T, 3) = P(T) P(T, 3) = 1/6; therefore, events are dependent because P(T, 3) doesn't equal P(T) P(T, 3) = 1/3; therefore, events are dependent because P(T, 3) doesn't equal P(T) P(T, 3) = 1/2; therefore, events are independent because P(T, 3) = P(T)

OpenStudy (b87lar):

do you have any guess on your mind?

OpenStudy (anonymous):

This one seems to be really hard for me. I feel like it would be independent, but I am not sure whether it's 1/12 or 1/2.

OpenStudy (b87lar):

your intuition is correct, they are independent. Given that the roll is 3, it is still same odd of 1/2 to see the coin landing tail.

OpenStudy (b87lar):

the important word here is "given"

OpenStudy (b87lar):

if it was "and" (as in "roll a 3 and land tail") that would be that 1/12

OpenStudy (b87lar):

makes sense?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (b87lar):

yw